Lion Inc., has sales of $2676, total assets of $769, and a debt−equity ratio of 0.7. If its return on equity is 21%.

What is Lion’s Net Income?

(Round final answer to 2 decimal places. Do not round intermediate calculations).
Answer #1

Please provide rating..

given that-
Sales 2676
total assets 769
debt−equity 0.7
return on equity 21%
debt % = =0.7/(1+0.7)
debt % = 41.18%
Equity = (1-debt) 58.82%
Equity= 769*58.82%                 452.35
Return on equity = Net income/equity
Net income = Return on equity * equity
Net income= 425.35*21%                       94.99